OpenStudy (anonymous):

Please HELP me. :3 will give medal and maybe fan :) Multiply and simplify: -3x^2y^2 times x^4

hartnn (hartnn):

so, you need to simplify \(x^2 \times x^4\) use the exponent rule \(\huge x^m \times x^n = x^{m+n}\)

hartnn (hartnn):

-3 and y^2 will remain as is

OpenStudy (anonymous):

Oh I forgot this A. -3x^5y^6 B. -3x^6y^2 C. 9x^5y^2 D. -3x^5y^2

hartnn (hartnn):

\(\Large x^2 \times x^4 = x^{2+4} = ... ? \)